OpenStudy (anonymous):

Solve by unfolding: a0 = 2, and, for n > 0, an = 7an-1 + 3. Hint: This one will involve applying the geometric sum formula

OpenStudy (anonymous):

I got to 3 + 3(7) + 3(7^2) + 3(7^3)...+7^n x (2) but I can't seem to simplify....

OpenStudy (anonymous):

assuming you are correct, use \[1+7+7^2+...+7^n=\frac{7^{n+1}-1}{7-1}\]

OpenStudy (anonymous):

So I would then have with my formula... \[3(\frac{ 7^{n+1}-1 }{ 6}) + 2 (7^{n})\]

OpenStudy (anonymous):

I think it needs simplified some more...

OpenStudy (anonymous):

And are you sure it's not \[3 (\frac{ 7^{n}-1 }{ 6 })\]